### v1

# Gentle spaghettification in effective LQG dust collapse

Abstract
Effective dust collapse inspired by loop quantum gravity predicts two main features for general realistic initial profiles: a quantum gravitational bounce of the stellar core, when the energy density becomes planckian, and shell-crossing singularities, arising within almost a planckian time after the bounce. The aim of this work is to study the mathematical and physical features of the effective spacetime near these singularities, through the Jacobi deviation equation. The results show that the radial and angular deviations between dust particles remain finite at both the bounce and shell-crossing singularity, providing a physical justification for extending the spacetime beyond shell-crossing singularities through weak solutions.
I Introduction
Spherically symmetric stellar collapse in classical general relativity generally leads to the formation of crush singularities, also called central singularities, according to the Penrose singularity theorem Penrose:1964wq . Even though such singularities have been widely studied in the literature and their mathematical properties analyzed in great detail, the common belief is that they are a pathological feature of Einstein theory of gravity, rather than a physical prediction. It has also been found that such singularities are not the only ones that develop during gravitational spherically symmetric collapse for generic initial density profiles. If neighboring layers of the collapsing star move at significantly different speeds, large inhomogeneity gradients can develop in the stellar energy density profile, sometimes leading to the so-called shell-crossing singularities (SCS) Penrose:1999vj ; Szekeres:1995gy ; Hellaby:1985zz ; Lasky:2006hq . Much less (though still significant) effort has been devoted to studying SCS compared to crush singularities, and the reasons are mainly threefold: firstly, during the collapse of pressureless matter, Hellaby and Lake proved Hellaby:1985zz that one can always choose initial data that do not develop such singularities before the central singularity is reached, which, in turn, is generally unavoidable. Additionally, by examining the mathematical features of these singularities, one can show Szekeres:1995gy that curvature scalars diverge less strongly compared to those associated with the crush singularities. Moreover, the behavior of matter falling toward these singularities is much less dramatic: despite the tidal forces being divergent at SCS, the angular and radial displacement of test particles remain finite Szekeres:1995gy . Furthermore, since the radial tidal force is negatively divergent, the so-called spaghettification effect is absent near SCS.
The purpose of this work is to understand how this picture changes once quantum gravitational corrections are taken into account. Among the possible candidate theories of quantum gravity, loop quantum gravity is one of the best-developed approaches Rovelli:2004tv ; Thiemann:2007pyv . Although a quantum gravitational model of stellar collapse is not yet available from the full theory, effective models have developed in recent years, where quantum gravitational corrections to Einstein equations are inspired by features of the underlying fundamental theory. The effective model on which this work is based is inspired from effective loop quantum cosmology, where the classical big bang singularity is replaced by a quantum gravitational bounce Ashtekar:2006wn . The modifications to Einstein field equations are based on the so-called $\bar{\mu}+K$-loop quantization scheme, where holonomy corrections with physical length $\sqrt{\Delta}$ (the minimum area gap in loop quantum gravity, proportional to the Planck length) affect only the angular component of the extrinsic curvature, and inverse triad corrections are neglected. The resulting effective dynamics has been extensively studied in recent years, both in the dust Husain:2021ojz ; Han:2023wxg ; Lewandowski:2022zce ; Giesel:2023hys ; Fazzini:2023scu ; Husain:2022gwp ; Kelly:2020lec ; Munch:2020czs ; Munch:2021oqn ; Cafaro:2024vrw ; Fazzini:2023ova ; Cipriani:2024nhx ; Bobula:2023kbo ; Bobula:2024chr ; Giesel:2023tsj and pressure cases Wilson-Ewing:2024uad ; Cafaro:2024lre . Despite much effort being devoted within this scheme to the Oppenheimer-Snyder collapse, the general picture regarding initial continuous decreasing inhomogeneous profiles is the following Fazzini:2023ova : the collapsing stellar core undergoes a quantum gravitational bounce when its energy density becomes planckian. Immediately after that, shell-crossing singularities arise, and the equations of motion break down. Although it is outside the scope of this work, for completeness, we mention that one can extend the dynamics beyond SCS formation by examining the integral form of the equation of motion, as in Painlevè-Gullstrand coordinates, it becomes a non-linear hyperbolic conservation law Leveque:2002 ; Batchelor_2000 , leading to the shockwave dynamics Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf . The same technique has been applied in the classical context Nolan:2003wp .
The question this work aims to address is how tidal forces behave during this effective dynamics, focusing in particular on the bounce point and shell-crossing singularities. The main results we found are the following: angular tidal forces remain bounded during the bounce of the stellar core and reach their maximum exactly at the bounce point. Unlike the classical case, there is a short period of time around the bounce point in which the angular tidal force is positive, and two zeroes are attained, before and after the bounce. The radial component of the tidal force, instead, is also bounded at the bouncing point, but diverges at SCS. However, unlike the classical case, where the tidal forces negatively diverges at SCS, in the effective case, the radial component of the tidal force is positive at SCS, meaning that spaghettification of a test body occurs, similarly to what happens near the classical central singularity. Despite this, importantly, by solving the radial component of the geodesic deviation equation, one finds that the radial separation between two points remains finite at the shell-crossing singularity, meaning that the spaghettification does not result in an infinite stretching of the body. This allows us to conclude that such singularities, arising during effective stellar collapse, should not be considered as dramatic as the central classical singularities, and it motivates the extension of spacetime beyond these singularities through weak solutions.
II Dust collapse in the marginally bound case
Effective dust collapse within the $\bar{\mu}+K-$loop quantization scheme has been extensively studied in Painlevè-Gullstrand coordinates Husain:2021ojz ; Kelly:2020lec ; Husain:2022gwp ; Fazzini:2023scu ; Cafaro:2024vrw ; Cipriani:2024nhx ; Fazzini:2025hsf and Lemaître-Tolman-Bondi coordinates (LTB) Giesel:2023hys ; Fazzini:2023ova ; Bobula:2024chr . The results derived in this work use LTB coordinates. The line element describing effective stellar collapse in these coordinates takes the following form
${\rm d}s^{2}=-{\rm d}t^{2}+\frac{(\partial_{R}r(R,t))^{2}}{1+\varepsilon(R)}{\rm d}R^{2}+r(R,t)^{2}{\rm d}\Omega^{2}~{},$ (1)
where $\varepsilon$ is a constant of motion fixed by the initial condition, and determines the Newtonian energy (kinetic plus potential) of any shell $R$ at infinity. We will focus here on the marginally bound case $\varepsilon(R)=0$, $\forall R$. The function $r(R,t)$ is the areal radius of each shell $R$ of the distribution at time $t$, and its explicit form is provided by the solution of the effective equations. In turn, the equation of motion in LTB coordinates takes the following form Giesel:2023hys :
$\left(\frac{\dot{r}}{r}\right)^{2}=\frac{F(R)}{r^{3}}\left(1-\frac{F(R)\gamma^{2}\Delta}{r^{3}}\right)~{},$ (2)
where we defined $F(R)\equiv 2Gm(R)$ for compactness, and $m(R)$ is the (conserved) gravitational mass within the $R-$shell. As one can explicitly notice from (2), the spacetime evolution is provided by the evolution of the areal radius of the shells composing the collapsing star. The analytic solution of the previous equation reads:
$r(R,t)=\left[F(R)\right]^{\frac{1}{3}}\left\{\frac{9}{4}\left[t-\alpha(R)\right]^{2}+\Delta\right\}^{\frac{1}{3}}~{},$ (3)
where $\alpha(R)$ is fixed by the initial condition. A fundamental feature of this generic solution, valid for any initial energy density profile, is that each shell $R$ will bounce at $t_{B}(R)=\alpha(R)$. In the pre-bounce (post-bounce) phase we have $t<\alpha(R)$ ($t>\alpha(R)$). Even though from (2) it seems that the shells’ dynamics is decoupled, and each shell evolves independently from the others (influenced only by the gravitational mass within its spherical volume), such solutions generally lead to shell-crossing singularities (SCS), at which points the equations cannot be trusted anymore. In order to define mathematically such singularities, we recall that the dust energy density in LTB coordinates reads
$\rho=\frac{\partial_{R}F}{8\pi Gr^{2}\partial_{R}r}~{}.$ (4)
A dynamical solution of the kind (3) develops SCS if, for a certain shell $R$, the following conditions are fulfilled: $\partial_{R}F\neq 0$, $\partial_{R}r(R,t)=0$. From (4), one can easily realize that these conditions imply a divergence in the dust energy density, and consequently a divergence in curvature scalars.
III Choice of initial profiles
The effective equations (2) hold for any initial energy density dust configuration in the marginally bound case. In particular, one can easily derive the effective cosmological dynamics Ashtekar:2006wn by requiring $\rho(t_{0},R)=const.$, as well as the vacuum solution Giesel:2023hys ; Giesel:2024mps , by imposing $F=const$. Here, we are interested in stellar collapse, so we will assume that the initial energy density profile is positive and of compact support. Moreover, to keep the result as general as possible, we will consider decreasing inhomogeneous profiles ($\partial_{R}\rho(t_{0},R)<0$) of compact support, without specifying the initial data.
For these kind of profiles, we can combine a theorem, stated and proved in Fazzini:2023ova , and two results, one from Fazzini:2023ova and the other from Fazzini:2025hsf :
Theorem.
For the case $\varepsilon=0$, a shell-crossing singularity forms if the initial distribution of the dust energy density $\rho(R,t_{0})$ is non-negative, continuous, of compact support, and for which m(R) is not everywhere zero,
and:
  1. 1.
    the latest time at which a shell-crossing singularity can arise in effective stellar collapse is $t=t_{B}(R)+(2/3)\sqrt{\Delta}$.
  2. 2.
    for initially decreasing continuous profiles, shell-crossing singularities cannot arise in a time $t\leq t_{B}(R)$.
Therefore, if we consider an initial profile that is non-negative, decreasing and of compact support, a shell-crossing singularity will necessarily form for a certain shell $R$ at time $t_{B}<t\leq t_{B}+(2\sqrt{\Delta}/3)$. These results allow us to restrict the analysis of tidal forces to the post-bounce dynamics of the shells, particularly to almost a planckian time interval after the bounce.
IV Geodesic deviation equation for LTB collapse
In order to understand tidal effects during generic dust collapse, we need to study the geodesic deviation equation. In the classical theory, given $u^{\mu}$ the 4-vector tangent to the geodesics, and $\delta x^{\nu}$, the displacement vector orthogonal to $u$ ($\delta x^{\nu}u_{\nu}=0$), the Jacobi equation holds Wald:1984rg
$\frac{D^{2}\delta x^{\mu}}{D\tau^{2}}=R^{\mu}_{\nu\rho\sigma}u^{\nu}u^{\rho}\delta x^{\sigma}~{},$ (5)
up to linear order in the deviation and its derivative. The covariant derivative is computed along the direction individuated by $u^{\mu}$, and therefore $\tau$ is the proper time along the geodesics. This equation does not uses the Einstein field equations but comes from the geometrical definition of the Riemann tensor. Therefore, it holds also at the effective level.
We are interested in the evolution of the deviation between dust particles during the stellar dynamics, and since dust particles follow geodesics also at the effective level (Appendix A), (5) describes tidal forces experienced by the dust itself. In particular, we are looking at a congruence of timelike geodesics for initially free-falling dust particles. Since the fluid $4$-velocity coincides with the $4$-velocity of a radially free-falling test particle, which in LTB coordinates takes the simple form $u^{\mu}=\{1,0,0,0\}$, equation (5) simplifies considerably
$\frac{D^{2}\delta x^{\mu}}{D\tau^{2}}=R^{\mu}_{00\sigma}\delta x^{\sigma}~{}.$ (6)
In order to get rid of the covariant derivative and simplify the equation further, we use a common procedure and introduce tetrads dInverno:1992gxs . The tetrads associated with the LTB line element (1), which coincide with the tetrad basis in the free-falling frame, in the marginally bound case $\varepsilon(R)=0$ take the following (classical) form:
$e^{\mu}_{0}=\left\{1,0,0,0\right\}$ (7)
$e^{\mu}_{1}=\left\{0,({\partial_{R}r})^{-1},0,0\right\}$ (8)
$e^{\mu}_{2}=\left\{0,0,r^{-1},0\right\}$ (9)
$e^{\mu}_{3}=\left\{0,0,0,({r\sin\theta})^{-1}\right\}~{}.$ (10)
These tetrads can be easily derived from $g^{\mu\nu}=e^{\mu}_{i}e^{\nu}_{j}\eta^{ij}$. Then, we write all the tensors appearing in (6) in terms of tetrads:
$R^{\mu}_{\nu\rho\sigma}=R^{a}_{bcd}e^{\mu}_{a}e^{b}_{\nu}e^{c}_{\rho}e^{d}_{\sigma}~{},$ (11)
$\delta x^{\mu}=e^{\mu}_{a}\delta x^{a}~{},$ (12)
where $a,b,c,d$ run from 00 to $3$. Notice that $\delta x^{0}=0$, since $u^{\mu}\delta x_{\mu}=0$ and $u^{\mu}$ has no spatial components. The deviation equation becomes
$\frac{D^{2}\delta x^{a}}{D\tau^{2}}=R^{a}_{00b}\delta x^{b}~{},$ (13)
which $0-$component is trivially satisfied. To arrive at (13) we used the tetrad postulate, which can be written as $De^{a}_{\mu}/D\tau=0$. The great advantage of using tetrads in LTB coordinates is that they define a locally inertial freely falling frame, for which the covariant derivative reduces to a simple ordinary derivative Szekeres:1995gy ; Fuchs(1990) . This leads to
$\frac{{\rm d}^{2}\delta x^{a}}{{\rm d}\tau^{2}}=R^{a}_{00b}\delta x^{b}~{}.$ (14)
To conclude, since the LTB coordinate time is the proper time of free-falling timelike observers, we can set $\tau=t$, and the equation further simplifies:
$\frac{\partial^{2}\delta x^{a}}{\partial t^{2}}=R^{a}_{00b}\delta x^{b}~{}.$ (15)
Although the explicit form of the effective Riemann tensor differs from the classical one, the effective LTB metric does not acquire quantum correction in its functional dependence from the solution $r(R,t)$, and the same holds for the Riemann tensor. Therefore, we can use the classical result obtained in Joshi:2024djy to write down (15) in a more explicit form:
$\partial^{2}_{t}{\delta x^{1}}=\frac{\partial^{2}_{t}\partial_{R}{r}}{\partial_{R}r}\delta x^{1}~{},$ (16)
${\partial^{2}_{t}\delta x^{i}}=\frac{\partial^{2}_{t}{r}}{r}\delta x^{i}~{},$ (17)
where the first equation determines the radial deviation and the index $i=2,3$ indicates the angular deviations in the tetradic basis. The time component of equation (15) gives a trivial identity. It is worth mentioning that, even though the functional dependence on $r$ takes the classical form, since the effective solution differs from the classical one, the same clearly holds for the solutions of equations (16) and (17).
V Tidal forces during effective stellar collapse
As pointed out before, effective inhomogeneous dust collapse does not end in the classical central singularity, which is replaced by a quantum gravitational bounce. In classical Einstein theory, angular tidal forces close to shell-crossing singularities do not diverge, while despite radial tidal forces being divergent, radial deviation does not go to zero at SCS, nor does it diverge Szekeres:1995gy . This makes such singularities less pathological than the central one. Here, we show that similar results hold also at the effective level.
To prove this, we compute explicitly (16), (17), using the solution (3) of the effective equation (2). Let’s start from the equation for angular deviation. After some computation, we find
$\partial^{2}_{t}\delta x^{i}=\frac{3\left[\Delta-\frac{3}{4}(t-\alpha)^{2}\right]}{2\left[\frac{9}{4}(t-\alpha)^{2}+\Delta\right]^{2}}\delta x^{i}~{}.$ (18)
Firstly, we notice that in the classical case $\Delta\rightarrow 0$, the angular tidal forces negatively diverges for $t\rightarrow\alpha(R)$ (classical central singularity). In contrast, here it remains always bounded, reaches its maximum at $t=\alpha(R)$ (the bouncing point), and two zeroes, for $t=\alpha\pm 2\sqrt{\Delta}/\sqrt{3}$.
The departure from the classical focusing behavior is clearly due to quantum gravity repulsion, which weakens the angular tidal force and, for almost a planckian time, produces a tidal angular repulsion.
Figure 1: Angular tidal force for $\alpha=1$, in the classical (dark blue curve) and effective collapse (red curve). In the classical case the function diverges at the crush singularity $t=\alpha=1$, while remains bounded at the effective level.
Fig.1 shows the comparison between the classical and effective cases. Notice that angular tidal forces do not diverge at shell-crossing singularity, which arises within a time $\alpha(R)<t\leq\alpha(R)+2\sqrt{\Delta}/3$.
It is worth mentioning that even though the angular tidal force remains bounded, it reaches a planckian value at the bounce, given by
$\partial^{2}_{t}\delta x^{i}|_{bounce}=\frac{3}{2\Delta}\delta x^{i}~{}.$ (19)
Let’s now examine the radial tidal force. By explicitly writing equation (16), after a lengthy but straightforward computation, we find
$\frac{\partial^{2}_{t}\partial_{R}r}{\partial_{R}r}=\left(\frac{F}{r^{3}}-\frac{10\Delta F^{2}}{r^{6}}\right)+\frac{\partial_{R}F}{\partial_{R}r}\left(\frac{4\Delta F}{r^{5}}-\frac{1}{2r^{2}}\right)~{}.$ (20)
Firstly, we notice that divergence can only occur if $\partial_{R}r=0$, provided the numerator of the last term is nonzero. This indicates that the radial tidal force remains finite at the bounce. To study the radial tidal force near and at SCS, we can assume that the first term is negligible
$\frac{\partial^{2}_{t}\partial_{R}r}{\partial_{R}r}\sim\frac{\partial_{R}F}{2r^{5}\partial_{R}r}\left(4\Delta F-r^{3}\right)~{}.$ (21)
Next, we recall that for the profiles we are considering (decreasing, continuous and of compact support), shell-crossing singularities arise for $\alpha<t\leq\alpha+2\sqrt{\Delta}/3$, which implies $F\Delta<r^{3}\leq 2F\Delta$, from (3). This means that the numerator of (21) is strictly positive, and at shell-crossing singularity the radial tidal force diverges. It is worth noting that such a divergence arises in the deep quantum gravitational regime and is a positive divergence: an infinite stretching radial force arises, conceptually similar to the one occurring classically near the central singularity, producing the spaghettification effect. In contrast, near classical shell-crossing singularities, the radial tidal force is negatively divergent: two points radially separated tend to be crushed together by the tidal force, even though in a weak fashion Szekeres:1995gy , as one can see by setting $\Delta=0$ in (21).
Let us now focus on the (positive) divergent behavior of the radial tidal force near and at the SCS. The condition for SCS can be rewritten as Fazzini:2023ova
$(t-\alpha)^{2}\partial_{R}m-2m(t-\alpha)\partial_{R}\alpha+\frac{4\Delta}{9}\partial_{R}m=0~{}.$ (22)
This is a second-degree equation in $t-\alpha$, with solutions $t_{1}=\alpha(R)+\delta t_{1}(R)$, $t_{2}=\alpha(R)+\delta t_{2}(R)$, with $\delta t_{1}\leq\delta t_{2}$. These solutions are guaranteed to be real for continuous, inhomogeneous, decreasing profiles of compact support, as shown in the previous section. It can be also easily shown that the solutions are distinct Fazzini:2023ova , except for an initial condition that satisfies $\frac{m\partial_{R}\alpha}{\partial_{R}m}=\frac{2\sqrt{\Delta}}{3}$, which is a fine-tuned case we do not consider here. Therefore, we can rewrite the condition for SCS as
$(t-\alpha(R)-\delta t_{1}(R))(t-\alpha(R)-\delta t_{2}(R))=0~{}.$ (23)
This means that close to $t_{SCS}=\alpha+\delta t_{1}$, the function $r^{\prime}$ goes to zero linearly: $r^{\prime}\sim K(t_{SCS},R)(t_{SCS}-t)$ for $t\lesssim t_{SCS}$, with $K(t_{SCS},R)>0$. Here we are only interested in tidal forces before and at the shell-crossing singularity. Beyond this point, the equation of motion in differential form breaks down, and analytic results are not reliable. With this in hand, we can rewrite (16):
$\partial^{2}_{t}\delta x^{1}\sim\frac{B(t_{SCS},R)}{t_{SCS}-t}\delta x^{1}~{},$ (24)
where $B(t_{SCS},R)$ is given by $\partial^{2}_{t}\partial_{R}r$ evaluated at shell-crossing singularity. This approximation holds only close to the SCS ($t\lesssim\alpha(R)+\delta t_{1}(R)$). Equation (24) can be solved analytically. Let us define $\delta t\equiv t_{SCS}-t>0$. The solution is
$\delta x^{1}(\delta{t})\sim$ $c_{1}\sqrt{\delta{t}}~{}I_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]+$ (25)
$c_{2}\sqrt{\delta{t}}~{}K_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]~{},$ (26)
where $I_{1}$ and $K_{1}$ are modified Bessel functions of the first and second kinds, and $c_{1},c_{2}$ are constants determined by initial conditions. The previous solution holds only for $\delta{t}\rightarrow 0^{+}$. While the first term goes to zero in this limit, the second term approaches a constant non-zero value. Their explicit Taylor expansions are:
$\sqrt{\delta{t}}I_{1}\left[2\sqrt{B(t_{SCS},R)\delta{t}}\right]=\sqrt{B(t_{SCS},R)}\delta{t}+O(\delta{t}^{2})~{},$ (27)
$\sqrt{\delta{t}}K_{1}\left[2\sqrt{B(R)\delta{t}}\right]=\frac{1}{2\sqrt{B(t_{SCS},R)}}+O(\delta{t})~{}.$ (28)
Therefore, even though the radial tidal forces diverge at SCS, two neighboring points that radially separated (and thus belong to two different matter layers) do not clash in general when shell-crossing singularity occur. For instance, these points can be thought of as belonging to two shells close to the ones that cross, but with a larger coordinate separation $\delta R$). If these points belong exactly to the crossing shells, they will clearly be crushed, and this condition is expected to be given by $c_{2}=0$ Szekeres:1995gy . This implies that the radial portions of the collapsing star (or even geodesic test particles radially separated) undergo spaghettification, but without being infinitely stretched. We refer to this effect as gentle spaghettification.
This result leads to the conclusion that SCS, which generally arise in effective stellar collapse, are less pathological than classical crush singularities. It also justifies the extension of the spacetime beyond SCS Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf , as done in the classical case through weak solutions. It is important to emphasize that this result is non-trivial. As mentioned before, even though the structure of the deviation equation follows the classical form, the effective solution differs from the classical one. Therefore, the behavior of tidal forces at shell-crossing singularities also differs from the classical behavior in the effective model.
To conclude, for profiles that develop shell-crossing singularities in the pre-bounce phase (where classical Einstein equations hold), the classical analysis Szekeres:1995gy can be directly exported to the effective model.
VI conclusions
Shell-crossing singularities are generally unavoidable in effective dust collapse inspired by loop quantum gravity, within the $\bar{\mu}+K$-loop quantization scheme. They therefore play a central role in such effective models, and their geometrical and physical features need to be investigated. In this work, we focused on the analysis of the tidal forces arising from effective dust collapse, with particular attention to the bouncing point and shell-crossing singularities. We found that tidal forces remain bounded at the bounce, in contrast to what happens in the classical case at the central crush singularity, and the same holds for the angular component at the shell-crossing singularity. However, the radial component of the tidal force positively diverges at the shell-crossing singularity.
Nonetheless, when studying the behavior of the radial deviation by solving the Jacobi equation, we find that it remains bounded during the whole dynamics, also when the shell-crossing singularity arises. The positive divergent behavior of the tidal force implies a spaghettification effect, similar to the one occurring near the central singularity in classical stellar collapse. However, the bounded behavior of the radial displacement between nearby points makes such an effect much less dramatic. We refer to this phenomenon as gentle spaghettification, and it allows the dynamics to extend beyond shell-crossing singularities in a meaningful way, through the integral form of the equations of motion. Future work in this direction will aim to investigate tidal forces in the effective gravitational collapse of dust in the non-marginally bound case Cipriani:2024nhx , and of fluids with pressure Cafaro:2024lre , in order to understand, at least at the numerical level, the role of pressure in the tidal forces experienced by test particles.
Acknowledgements.
This work is supported in part by the Natural Sciences and Engineering Research Council of Canada.
Appendix A Effective geodesic flow of dust particles
At the effective level the energy-momentum tensor is covariantly conserved
$\nabla^{\mu}T_{\mu\nu}=0~{}.$ (29)
This is a direct consequence of the fact that the matter part of the Einstein equation is not modified by quantum corrections. The energy-momentum tensor for dust reads $T_{\mu\nu}=\rho u_{\mu}u_{\nu}$, where $u^{\mu}$ is the fluid $4$-velocity in its rest frame. By imposing (29) one can easily get $\rho u^{\mu}\nabla_{\mu}u^{\nu}=0$, which for $\rho\neq 0$ implies that dust particles follow geodesics also at the effective level.